Explain the significanceof azimuthal quantum numbers. Indicate the possible values of l and m1 for n=3
1
Expert's answer
2017-09-21T03:24:07-0400
The azimuthal quantum number is a quantum number for an atomic orbital that determines its orbital angular momentum and describes the shape of the orbital. The azimuthal quantum number is the second of a set of quantum numbers which describe the unique quantum state of an electron (the others being the principal quantum number, following spectroscopic notation, the magnetic quantum number, and the spin quantum number). It is also known as the orbital angular momentum quantum number, orbital quantum number or second quantum number, and is symbolized as ℓ.

l = n-1 = 3 - 1 = 2
m = -2 , -1 , 0 , 1 , 2